calculate operating surplus and net exports

wages and salaries 2400

domestic income 4000

gross domestic fixed capital formation 1000

mixed income of self employed 400

private final consumption expenditure 2000

net factor income from abroad 200

net indirect taxes 150

government final consumption expenditure 1000

consumption of fixed capital 100

profit 500

change in stock 100


1
Expert's answer
2021-05-07T09:54:35-0400

operating surplus, 

Operating surplus=NDPFc−wages and salaries−mixed income from self employed


The net domestic product (factor cost), 

NDPFc=GDPMp(domestic income−consumption of fixed capital(depreciation−Net indirect

taxes

NDPFc"=4000\u2212100\u2212150"

NDPFc"=3750"  

Substitute the values in the formula, 

Operating surplus"=3750\u22122400\u2212400"

Operating surplus"=950"

But we add profit in this too. 

Operating surplus"=950+\space profit"

Operating surplus"=950+500"

Operating surplus"=1450"


GDPmp=private final consumption expenditure+gross domestic fixed capital formation

+government final consumption expenditure+net exports

Substitute the values in the formula,

"4000=2000+1000+1000+Net\space exports"

"4000=4000+Net \space exports"

Net exports"=0"

operating surplus is 1450 and net exports are 0.